I am not sure what I am supoose to do. My teacher gave us questions for a pre-quiz and this was one of the questions. I have no idea what i am suppose to do or how to convert inches to meters.
qin Determine L in meters
Answer #1

This looks like a right angle triangle. Therefore, you can apply the Pythagoras theorem. It states that in a right angle triangle, the square of the hypotenuse is equal to the sum of squares of the lengths of other two sides. Hence, in this case, since L is the hypotenuse:

L ^ 2 = (9 ^ 2) + (16 ^ 2)

L ^ 2 = 81 + 256

L ^ 2 = 337

L = 18.3576 inches.

Now 1 inch = 0.0254 meters

Therefore, L = 0.4663 m